a solution of 78% alcohol is to be mixed with a solution of 23% alcohol to form 165 liters of a 58% solution. How many liters of the 78% solution must be used?
Answer by Alan3354(69443)

I just did this one.
----------------
a solution of 78% alcohol is to be mixed with a solution of 23% alcohol to form 165 liters of a 58% solution. How many liters of the 78% solution must be used?
---------
s = amount of 78%
t = amount of 23%
-------
s + t = 165 (total solution)
--> t = 165 - s
-------
0.78t + 0.23s = 0.58*165 (total alcohol)
Sub for t
0.78s + 0.23(165-s) = 0.58*165
78s + 3795 - 23s = 9570
55s = 5775
s = 105 liters of 78%